OpenStudy (anonymous):

How do you do this the format is y-y1=m(X-X1) The problem: (-8,3), m= -3/5 whats the answer is point slope form?

OpenStudy (anonymous):

I NEED HELP PLEASE>

OpenStudy (anonymous):

which one is the answer?

jimthompson5910 (jim_thompson5910):

y - y1 = m(x - x1) y - y1 = -3/5(x - x1) ... plug in m = -3/5 y - 3 = -3/5(x - (-8)) ... plug in (-8, 3) y - 3 = -3/5(x + 8)

jimthompson5910 (jim_thompson5910):

the last line is the most simplified answer

jimthompson5910 (jim_thompson5910):

the other lines are steps leading up to the answer

OpenStudy (anonymous):

thank you so much :)

jimthompson5910 (jim_thompson5910):

np